### v1

# Hidden Bose-Einstein Singularities in Correlated Electron Systems

Abstract
Hidden singularities in correlated electron systems, which are caused by pair fluctuations of electron-electron or electron-hole bubbles obeying Bose-Einstein statistics, are unveiled theoretically. The correlation function of each pair fluctuation is shown to have a bound in the zero Matsubara frequency branch, similarly as the chemical potential of ideal Bose gases. Once the bound is reached, the self-energy starts to acquire a component proportional to Green’s function itself, i.e. the structure called one-particle reducible, to keep the correlation function within the bound. The singularities are closely related with, but distinct from, phase transitions with broken symmetries. Passing down through them necessarily accompanies a change in the single-particle density of states around the excitation threshold, such as the pseudo-gap behavior found here for the negative-$U$ Hubbard model above the superconducting transition temperature.
I Introduction
Systems of identical particles with integer spins embrace a unique phase transition called Bose-Einstein condensation characterized by the emergence of macroscopic quantum coherence Cornell95 ; Ketterle97 ; Leggett06 , which may be regarded as realization of the singularity of the Bose distribution function. Specifically for noninteracting systems below the critical temperature, the chemical potential is pinned to the value of the lowest energy level with the singularity, on which particles accumulate macroscopically. The purposes of the present paper are to show that similar singularities are also present in the pair-fluctuation channels of fermion systems described in terms of the Bose distribution function, and to clarify some basic features of the hidden phases realized after the singularities have been reached.
The singularities are relevant to logarithmic contributions to the grand thermodynamic potential given diagrammatically in terms of closed electron-electron or electron-hole bubbles; see Fig. 1 below. The argument of the logarithm can be expressed in terms of boson Matsubara frequencies, and the vanishing of it in the zero-frequency branch has been identified as the second-order transition point in the perturbative treatments of using the bare Green’s function, known as the Thouless criterion for superconductivity Thouless60 ; NSR85 and Stoner criterion for ferromagnetism Stoner38 ; Blundell01 , for example; they are also identical with the mean-field estimations of the critical temperatures as well as those based on the random-phase approximations IKK63 ; BS66 ; DE66 . Making Green’s function self-consistent and incorporating terms other than the logarithmic contributions, however, each singular point generally becomes different from the corresponding transition point with a broken symmetry, as shown below both analytically and numerically. It also has not been clarified what happens after the singular points are reached and how to describe the phases theoretically. These are the issues to be studied here.
This paper is organized as follows. In Sect. 2, we elaborate the hidden singularities and introduce a procedure to handle them theoretically. In Sect. 3, we present numerical results on the negative-$U$ Hubbard model. In Sect. 4, we provide concluding remarks. We use the units of $\hbar=k_{\rm B}=1$.
Figure 1: Diagrammatic expression of $\Phi$ in the FLEX-S approximation Kita11 . A small circle denotes the interaction vertex with particle-hole and spin degrees of freedom, and a line denotes an element of $\hat{G}$.
II Hidden Singularities
II.1 Luttinger-Ward functional
One of the theoretical frameworks suitable for the present purposes is the Luttinger-Ward expression of the grand thermodynamic potential $\Omega$ given as a functional of Green’s function $\hat{G}$ LW60 ; DDM64 ; DDM64-2 ; HRCZ07 ; Kita11 . The stationarity condition of $\Omega$ in terms of $\hat{G}$ yields a self-consistent equation for $\hat{G}$, which have been known by various names such as Dyson’s, the Schwinger-Dyson, or Dyson-Gor’kov equation. Being nonlinear in $\hat{G}$, it can describe ordered phases with spontaneous broken symmetries continuously from the normal state on the same footing. Indeed, the ordered phases are characterized by the emergence of some novel component in the matrix $\hat{G}$ with particle-hole and spin degrees of freedom, such as the anomalous Green’s function for superconductivity and spin polarization in $\hat{G}$ for ferromagnetism. This Luttinger-Ward functional with full particle-hole and spin degrees of freedom can be written as Kita11
$\Omega=-\frac{1}{2\beta}{\rm Tr}\,\bigl{[}\ln\bigl{(}-\hat{G}_{0}^{-1}+\hat{\Sigma}\bigr{)}+\hat{\Sigma}\hat{G}\bigr{]}+\Phi,$ (1)
where $\beta$ is the inverse temperature, $\hat{G}_{0}$ is the noninteracting Green’s function, $\hat{\Sigma}$ is the self-energy, and Tr denotes trace over the arguments of $\hat{G}$. The stationarity condition of Eq. (1) with respect to $\hat{G}$ yields the Schwinger-Dyson equation
[…]
These are formally the exact expressions, and practical calculations can be performed based on some approximate expression for $\Phi\!=\!\Phi[\hat{G}]$, obeying Eq. (2) just like the exact theory. The scheme of Eq. (2) has also been known as conserving approximations of Kadanoff and Baym KB62 or $\Phi$-derivable approximations Baym62 . The potential $\Omega$ as a functional of $\hat{G}$ is alternatively called effective action in relativistic quantum field theory CJT74 ; Weinberg95 .
II.2 Normal state
To elucidate the hidden singularities that lies near phase transitions with broken symmetries, we specifically consider the negative-$U$ Hubbard model at a low electron density $\bar{n}\!=\!N/V$, which has negligible lattice anisotropy and definitely exhibit $s$-wave superconductivity at low temperatures. We study this system based on the fluctuation-exchange approximation for superconductivity (FLEX-S approximation), whose $\Phi$ is given diagrammatically as Fig. 1 Kita11 . We first focus on the normal state, where the series reduces to the $\Phi$ functional of the standard fluctuation-exchange (FLEX) approximation BSW89 ; BS89 ; Kita11 ,
$\Phi_{\rm n}=$ $\,\frac{UN^{2}}{4V}+\frac{1}{\beta}\sum_{\vec{q}}\Bigl{[}\ln(1-x)+x\Bigr{]}_{x=U\chi_{G\bar{G}}(\vec{q}\,)}$
$\,+\frac{3}{2\beta}\sum_{\vec{q}}\biggl{[}\ln(1-x)+x+\frac{1}{2}x^{2}\biggr{]}_{x=U\chi_{GG}(\vec{q}\,)}$
$\,+\frac{1}{2\beta}\sum_{\vec{q}}\biggl{[}\ln(1+x)-x+\frac{1}{2}x^{2}\biggr{]}_{x=U\chi_{GG}(\vec{q}\,)}\,.$ (3)
Here, $\vec{q}\equiv({\bf q},i\omega_{\ell})$ is the four momentum consisting of wavevector ${\bf q}$ and boson Matsubara frequency $\omega_{\ell}$, and functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ are defined in terms of the normal-state Green’s function $G(\vec{k})$ and its time-reversed partner $\bar{G}(\vec{k})\!\equiv\!G(-\vec{k})$ generally by
$\chi_{AB}(\vec{q}\,)\equiv$ $\,-\frac{1}{\beta V}\sum_{\vec{k}}A(\vec{k}+\vec{q}\,)B(\vec{k}),$ (4a)
where $\vec{k}\!\equiv\!({\bf k},i\varepsilon_{n})$ with $\varepsilon_{n}$ denoting the fermion Matsubara frequency; this function satisfies
$\chi_{AB}(\vec{q}\,)\!=\!\chi_{BA}(-\vec{q}\,)\!=\!\chi_{\bar{A}\bar{B}}(-\vec{q}\,).$ (4b)
as can be shown by a change of summation variable $\vec{k}$. The first term in Eq. (3) is the Hartree-Fock contribution, the second one is the particle-particle ladder series relevant to the superconducting fluctuations including the second-order contribution, and the third and fourth terms represent the spin and charge fluctuations, respectively, composed of the particle-hole ladder and ring diagrams BSW89 . The corresponding self-energy is obtained by the functional differentiation of Eq. (2); for the normal state that is symmetric in the spin and particle-hole spaces, the differentiation simplifies to $\Sigma_{\rm n}(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi_{n}}{\delta G(\vec{k})}$ in terms of Eq. (3), yielding BSW89 ; BS89 ; Kita11 ; Yanase03 ; Kontani13
$\Sigma_{\rm n}(\vec{k})=$ $\,\frac{UN}{2V}+\frac{U}{\beta V}\sum_{\vec{q}}\frac{U\chi_{G\bar{G}}(\vec{q}\,)}{1-U\chi_{G\bar{G}}(\vec{q}\,)}G(-\vec{k}+\vec{q}\,)$
$\,+\frac{3U}{2\beta V}\sum_{\vec{q}}\frac{\bigl{[}U\chi_{GG}(\vec{q}\,)\bigr{]}^{2}}{1-U\chi_{GG}(\vec{q}\,)}G(\vec{k}-\vec{q}\,)$
$\,-\frac{U}{2\beta V}\sum_{\vec{q}}\frac{\bigl{[}U\chi_{GG}(\vec{q}\,)\bigr{]}^{2}}{1+U\chi_{GG}(\vec{q}\,)}G(\vec{k}-\vec{q}\,).$ (5)
Now, we are ready to discuss the hidden singularities in full detail. With replacing $G$ by $G_{0}$ in Eq. (5), the conditions of the vanishing denominators at $\vec{q}=0$ in the second and third terms reproduce the Thouless criterion $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ for superconductivity NSR85 and Stoner criterion $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ for ferromagnetism Blundell01 , respectively. However, the renormalized conditions $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ do not specify the phase transition points of respective broken symmetries. Indeed, by linearizing the self-consistent equation for the anomalous Green’s function $F$ in the FLEX-S approximation, we obtain the equation for the superconducting transition temperature $T_{\rm c}$ that the eigenvalue problem
$\sum_{\vec{k}^{\prime}}\left[\delta_{\vec{k}\vec{k}^{\prime}}+\frac{1}{\beta V}U^{\rm s}_{\vec{k}\vec{k}^{\prime}}G(\vec{k}^{\prime})G(-\vec{k}^{\prime})\right]\Delta(\vec{k}^{\prime})=0$ (6)
have a non-trivial anomalous self-energy $\Delta(\vec{k})$ due to the emergence of a zero eigenvalue, where $U^{\rm s}_{\vec{k}\vec{k}^{\prime}}$ is the pairing interaction with correlations given explicitly by Eq. (91) of Ref. Kita11, , i.e.,
$U_{\vec{k}\vec{k}^{\prime}}^{\rm s}=$ $\,U\!\left[1-2x+\frac{3}{2}\frac{x}{1-x}-\frac{1}{2}\frac{x}{1+x}\right]_{x=U\chi_{GG}(\vec{k}-\vec{k}^{\prime})}$
$\,-\frac{U^{2}}{\beta V}\sum_{\vec{q}}\frac{(G_{\vec{k}+\vec{q}}+G_{-\vec{k}+\vec{q}})G_{\vec{k}^{\prime}+\vec{q}}}{\bigl{[}1+U\chi_{GG}(\vec{q})\bigr{]}\bigl{[}1-U\chi_{G\bar{G}}(\vec{q})\bigr{]}}.$ (7)
This condition for $T_{\rm c}$, which is also identical with the one that the superconducting correlation function diverges, is manifestly different from $U\chi_{G\bar{G}}(\vec{0})\!=\!1$. The same statement holds true for the case of ferromagnetism with $U\!>\!0$, as can be checked easily based on the consideration of incorporating the spin polarization in Green’s function. Hence, we should allow for the possibility that $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ are reached independently of the phase transitions with broken symmetries. The fact that $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ and $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ specify the mean-field $T_{\rm c}$ suggests that the renormalized ones will be realized near the phase transition points with broken symmetries.
Another key observation here is that the values of $U\chi_{G\bar{G}}(\vec{0})$ and $U\chi_{GG}(\vec{0})$ cannot exceed 1 physically. This can be seen as follows. The zero Matsubara frequency contribution to the second term of Eq. (12) can be written as
$\frac{1}{\beta}\sum_{\bf q}\left\{\ln\left[1-U\chi_{G\bar{G}}(q,0)\right]+U\chi_{G\bar{G}}(q,0)\right\},$ (8)
where $U\chi_{G\bar{G}}(q,0)$ is real, as can be shown based on Eq. (4). Since Eq. (8) is part of the grand thermodynamic potential which is real, the inequality $11-U\chi_{G\bar{G}}(q,0)\!\geq\!$ 1 should hold. Put it another way, the series expansion of Eq. (8) does not converge for $U\chi_{G\bar{G}}(q,0)>1$. Thus, we are led to the conclusion that $U\chi_{G\bar{G}}(\vec{0})$ and $U\chi_{GG}(\vec{0})$ must have a common bound 1, at which they should be pinned after reaching it; we call $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ hidden singularities.
The singularities can be defined more generally beyond the FLEX approximation in terms of the logarithmic series of Fig. 1 with the bare interaction $U$ replaced by some renormalized one $\underline{\Gamma}$: the matrix argument of the logarithm $\ln(\underline{1}+\underline{\Gamma}\,\underline{GG})$ comes to have a zero eigenvalue in the zero Matsubara frequency branch.Kita11 ; DDM64-2 Thus, $U$ in our analysis may be regarded as a renormalized effective interaction.
[…]
We propose a resolution to this issue within the variational framework of Eqs. (1) and Eq. (2) by making use of the method of the Lagrange multiplier so as not to affect $\Omega$ explicitly. Specifically for the negative-$U$ case where $U\chi_{G\bar{G}}(\vec{0})$ is relevant, we add to the $\Phi$ functional the following term that vanishes for $U\chi_{G\bar{G}}(\vec{0})=1$,
$\varDelta\Phi_{\lambda}^{\rm n}\equiv\lambda_{\rm n}N\left[1-U\chi_{G\bar{G}}(\vec{0})\right],$ (9)
where the particle number $N$ is incorporated in the prefactor to make the Lagrange multiplier $\lambda_{\rm n}$ intensive; note that $\lambda_{\rm n}$ has the dimension of energy. Performing the differentiation of Eq. (2) with Eq. (9) added to $\Phi$, we obtain the expression of the self-energy for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})=1$ in terms of Eq. (5) as
$\Sigma_{\lambda}^{\rm n}(\vec{k})=\Sigma_{\rm n}(\vec{k})+\lambda_{\rm n}N\frac{U}{V}G(-\vec{k}).$ (10)
The additional second term stemming from Eq. (9) is characteristic of the pinned phase, which corresponds to the condensate fraction of the ideal Bose-Einstein condensation Leggett06 . It has the unusual structure called one-particle reducible (1PR), i.e. the structure that becomes disconnected upon cutting a line, which originates from the two-particle reducible structure of Eq. (9). The parameter $\lambda_{\rm n}$ in Eq. (9) enables us to control the value of $U\chi_{G\bar{G}}(\vec{0})$ and should be chosen in such a way as to meet $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ by solving Dyson’s equation
$G(\vec{k})\!=\!\bigl{[}G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}^{\rm n}(\vec{k})\bigr{]}^{-1}$ (11)
self-consistently. This completes the formulation for the normal state.
The parameter $\lambda_{\rm n}$ is finite only for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and expected to grow almost linearly as a function of temperature initially. This linear change in the structure of the self-energy need not be considered in performing the first derivatives of $\Omega$ because of the stationarity condition (2). However, it does bring singularities in the second derivatives of $\Omega$ such as the specific heat, similarly as those caused by the emergences of $\Delta(\vec{k})$ for superconductivity and the spin-polarized self-energy for ferromagnetism near $T_{\rm c}$. Thus, the point $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ may be regarded as specifying a kind of transition without apparent broken symmetries. Detailed studies on the observable singularities remain to be performed in the future.
II.3 Superconducting state
The above consideration for the normal state can be extended directly and continuously to the $s$-wave superconducting state characterized by the emergence of the anomalous Green’s function $F(\vec{k})\!=\!F(-\vec{k})$. Specifically, we can construct the $\Phi$ functional in the FLEX-S approximation so as to incorporate all the anomalous processes derivable from those of Eq. (3) based on the diagrams of Fig. 1 Kita11 ,
$\Phi=$ $\,-\frac{U}{\beta}\sum_{\vec{q}}\chi_{+}(\vec{q}\,)-\frac{1}{2\beta}\sum_{\vec{q}}\bigl{[}U\chi_{+}(\vec{q}\,)\bigr{]}^{2}$
$\,+\frac{1}{2\beta}\sum_{\vec{q}}{\rm Tr}\,\biggl{[}\ln\bigl{(}\underline{1}+\underline{x}\bigr{)}-\underline{x}+\frac{1}{2}\underline{x}^{2}\biggr{]}_{\underline{x}=U\underline{\chi}^{({\rm c})}(\vec{q}\,)}$
$\,+\frac{3}{2\beta}\sum_{\vec{q}}\biggl{[}\ln\bigl{(}1-x\bigr{)}+x+\frac{1}{2}x^{2}\biggr{]}_{x=U\chi_{+}(\vec{q}\,)}.$ (12)
Here $\underline{1}$ is the $3\times 3$ unit matrix, $\underline{\chi}^{({\rm c})}(\vec{q}\,)$ is defined in terms of Eq. (4) with $A,B\!=\!G,\bar{G},F$ by
$\underline{\chi}^{({\rm c})}(\vec{q}\,)\equiv\begin{bmatrix}\vspace{1mm}\chi_{-}(\vec{q}\,)&\sqrt{2}\chi_{GF}(\vec{q}\,)&-\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)\\ \vspace{1mm}\sqrt{2}\chi_{GF}(\vec{q}\,)&-\chi_{G\bar{G}}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)\\ -\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)&-\chi_{\bar{G}G}(\vec{q}\,)\end{bmatrix},$ (13a)
and $\chi_{\pm}(\vec{q}\,)$ denote
$\chi_{\pm}(\vec{q}\,)\equiv$ $\,\chi_{GG}(\vec{q}\,)\pm\chi_{FF}(\vec{q}\,).$ (13b)
Matrix $\underline{\chi}^{({\rm c})}(\vec{q}\,)$ represents mixing of the superconducting and charge fluctuations due to the emergence of $F$. Equation (9) is now replaced by
$\varDelta\Phi_{\lambda}=\lambda N\det\bigl{[}\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)\bigr{]},$ (14)
where $\lambda$ is determined by the condition that the smallest eigenvalue of the matrix $\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)$ stay zero once the value has been reached. Although it should be removed when all the eigenvalues are positive, our numerical study given below in Sect. III presents an example that incorporating $\varDelta\Phi_{\lambda}$ is indispensable to obtaining physical (i.e. finite) solutions throughout the superconducting phase.
The contributions of Eq. (12) to the normal and anomalous self-energies can be calculated by $\Sigma(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta G(\vec{k})}$ and $\Delta(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi}{\delta F(\vec{k})}$, yielding Kita11
$\Sigma(\vec{k})=$ $\,\frac{UN}{2V}+\frac{1}{\beta V}\sum_{\vec{q}}U_{22}^{\rm eff}(\vec{q}\,)G(-\vec{k}+\vec{q}\,)$
$\,+\frac{1}{2\beta V}\sum_{\vec{q}}\bigl{[}3U_{+}^{\rm eff}(\vec{q}\,)-U_{11}^{\rm eff}(\vec{q}\,)\bigr{]}G(\vec{k}-\vec{q}\,)$
$\,-\frac{\sqrt{2}}{\beta V}\sum_{\vec{q}}U_{12}^{\rm eff}(\vec{q}\,)F(\vec{k}-\vec{q}\,),$ (15a)
$\Delta(\vec{k})=$ $\,\frac{U}{\beta V}\sum_{\vec{k}^{\prime}}F(\vec{k}^{\prime})+\frac{1}{\beta V}\sum_{\vec{q}}U_{23}^{\rm eff}(\vec{q}\,)F(\vec{k}-\vec{q}\,)$
$\,+\frac{1}{2\beta V}\sum_{\vec{q}}\bigl{[}3U_{+}^{\rm eff}(\vec{q}\,)+U_{11}^{\rm eff}(\vec{q}\,)\bigr{]}F(\vec{k}-\vec{q}\,)$
$\,-\frac{1}{\beta V}\sum_{\vec{q}}U_{12}^{\rm eff}(\vec{q}\,)\frac{G(\vec{k}+\vec{q}\,)+G(-\vec{k}+\vec{q}\,)}{\sqrt{2}}.$ (15b)
Here $U_{+}^{\rm eff}(\vec{q}\,)$ is defined by
$U_{+}^{\rm eff}(\vec{q}\,)\equiv U\biggl{[}\frac{U\chi_{+}(\vec{q}\,)}{1-U\chi_{+}(\vec{q}\,)}-\frac{1}{3}U\chi_{+}(\vec{q}\,)\biggr{]},$ (16a)
and $U_{ij}^{\rm eff}(\vec{q}\,)$ denotes the $ij$ element of the matrix
$\underline{U}^{\rm eff}(\vec{q}\,)\equiv U\bigl{[}U\underline{\chi}^{({\rm c})}(\vec{q}\,)\bigr{]}^{2}\bigl{[}\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)\bigr{]}^{-1}.$ (16b)
After adding Eq. (14) to the $\Phi$ functional, the self-energies acquire terms proportional to Green’s functions as
$\Sigma_{\lambda}(\vec{k})=$ $\,\Sigma(\vec{k})+U\frac{N}{V}\left[\lambda_{\bar{G}}G(-\vec{k})+\lambda_{G}G(\vec{k})+\lambda_{F}F(\vec{k})\right],$ (17a)
$\Delta_{\lambda}(\vec{k})=$ $\,\Delta(\vec{k})+U\frac{N}{V}\left\{\tilde{\lambda}_{F}F(\vec{k})+\tilde{\lambda}_{G}[G(\vec{k})+G(-\vec{k})]\right\},$ (17b)
with
$\lambda_{\bar{G}}\!\equiv\!2\lambda\bigl{\{}\bigl{[}1\!-\!U\chi_{\bar{G}G}(\vec{0})\bigr{]}\bigl{[}1\!+\!U\chi_{-}(\vec{0})\bigr{]}\!-\!2\bigl{[}U\chi_{\bar{G}F}(\vec{0})\bigr{]}^{2}\bigr{\}},$
for example. The Dyson-Gor’kov equation can be written in terms of these self-energies as
$\begin{bmatrix}G(\vec{k})\\ F(\vec{k})\end{bmatrix}=\frac{1}{D(\vec{k})}\begin{bmatrix}-G_{0}^{-1}(-\vec{k})+\Sigma_{\lambda}(-\vec{k})\\ \Delta_{\lambda}(\vec{k})\end{bmatrix},$ (18)
with $D(\vec{k})\!\equiv\![G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}(\vec{k})][-G_{0}^{-1}(-\vec{k})\!+\!\Sigma_{\lambda}(-\vec{k})]\!-\![\Delta_{\lambda}(\vec{k})]^{2}$. It has turned out numerically that $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ holds for $\lambda$ that meets $\det[\underline{1}+U\underline{\chi}^{({\rm c})}(\vec{q}\,)]=0$. Note that $\lambda_{\bar{G}}$ should reach $\lambda_{\rm n}$ of Eq. (10) at the superconducting transition temperature. This completes our formulation to study the superconducting phase continuously from the normal state based on the FLEX-S approximation.
III Numerical Results
Self-consistent calculations of Dyson’s and the Dyson-Gor’kov equations have been performed numerically on the negative-$U$ Hubbard model at a low electron density under the condition of the constant density
$\frac{N}{V}=\frac{2}{\beta V}\sum_{\vec{k}}G(\vec{k})\,e^{i\varepsilon_{n}0_{+}},$ (19)
where $0_{+}$ denotes an infinitesimal positive constant; the technical details are given in Appendix. We have adopted the units of $12m\!=\!\hbar\!=\!k_{\rm F}\!=\!$ 1, where $m$ is the electron mass and $k_{\rm F}$ is the Fermi momentum. The momentum cutoff $k_{\rm c}\!=\!10$ has been introduced to describe a finite band width. Specifically, our $G_{0}^{-1}(\vec{k})$ can be written as $G_{0}^{-1}(\vec{k})\!=\!i\varepsilon_{n}\!-\!k^{2}\!+\!\mu$ with $100\!\leq\!k\!\leq\!$ 10, where $\mu$ is the chemical potential determined by Eq. (19). We have chosen the parameter $U$ as $UN(0)=-0.09$ in terms of the density of states per spin and volume at the Fermi energy $N(0)\!=\!\frac{1}{4\pi^{2}}$; note also $\frac{N}{V}\!=\!\frac{1}{3\pi^{2}}$ in the present units. The corresponding mean-field superconducting transition temperature $T_{{\rm c}0}$, which is evaluated by using only the first terms on the right-hand sides of Eqs. (15a) and (15b), has the value $T_{{\rm c}0}\!=\!0.1767$ that is one order of magnitude smaller than the noninteracting Fermi energy $\varepsilon_{\rm F}^{0}\!=\!1$. Thus, our system lies in the intermediate coupling regime.
Figure 2: Temperature dependences of the Lagrange multipliers $(\lambda_{\rm n},\lambda_{\bar{G}})$ and the first-order anomalous self-energy $\Delta_{1}$ in comparison with that of the mean-field energy gap $\Delta_{\rm MF}$.
Figure 2 plots temperature dependences of the Lagrange multipliers $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ of Eqs. (10) and (17a), respectively, together with the first-order anomalous self-energy $\Delta_{1}$ defined by the first term on the right-hand side of Eq. (15b) in terms of fully self-consistent $F$ with correlations, which is to be compared with the mean-field energy gap $\Delta_{\rm MF}$. The crosses and squares denote numerical results, and smooth curves are obtained by interpolating them. Matching the curves of $\lambda_{\rm n}$ and $\lambda_{\bar{G}}$ yields an estimate of the superconducting transition temperature $T_{{\rm c}}\!=\!0.0654$ with correlations, which is far below the mean-field value $T_{{\rm c}0}\!=\!0.1767$. This mean-field calculation yields the ratio $\Delta_{\rm MF}(0)/T_{{\rm c}0}\!=\!1.635$ that is still close to the weak-coupling value $1.76$ Leggett06 , while the one with correlations evaluated by $\Delta_{1}(0)/T_{{\rm c}}$ is enhanced up to $2.68$. Note that all the quantities in Fig. 2, which originally have the dimension of energy, are of the same order as $|UN(0)|=0.09$, as may be expected naturally.
Figure 3: (a) Real and (b) imaginary parts of the correlation self-energy $\Sigma_{\rm c}(k,i\varepsilon_{n})$ calculated at $T\!=\!0.08$, which is expressed as a sum $\Sigma_{\rm c}\!=\!\Sigma_{\rm c}^{\rm i}\!+\!\Sigma_{\rm c}^{\rm r}$ of the 1PI and 1PR contributions $\Sigma_{\rm c}^{\rm i}$ and $\Sigma_{\rm c}^{\rm r}$, respectively.
One of the key features of Fig. 2 is that $\lambda_{\rm n}$ starts to develop well above $T_{\rm c}\!=\!0.0654$ at $T_{1{\rm PR}}\!=\!0.100$, which means that the singularity of $U\chi_{G\bar{G}}(0)\!=\!1$ is reached already in the normal state. This is a continuous transition with no apparent broken symmetries, which is characterized by the emergence of the 1PR structure in the self-energy. Figure 3 shows (a) real and (b) imaginary parts of the correlation self-energy $\Sigma_{{\rm c}}$, i.e., Eq. (10) without the $\frac{UN}{2V}$ term in Eq. (5), calculated at $T\!=\!0.08$, where $\Sigma_{{\rm c}}^{\rm i}$ and $\Sigma_{{\rm c}}^{\rm r}$ denote the one-particle-irreducible (1PI) and 1PR contributions, respectively. Thus, the emergence of $\Sigma_{{\rm c}}^{\rm r}$ brings a marked change in the structure of the self-energy, especially around the Fermi momentum at low Matsubara frequencies, as expected naturally from $\Sigma_{{\rm c}}^{\rm r}(\vec{k})\!\propto\!G(-\vec{k})$.
Figure 4: Anomalous self-energy $\Delta_{\lambda}(k,i\varepsilon_{n})$ calculated at $T\!=\!0.0353$.
The 1PR structure is found to persist below $T_{1{\rm PR}}$ into the superconducting phase and emerges also in the anomalous self-energy $\Delta_{\lambda}$. As seen in Fig. 4 calculated at $T\!=\!0.2T_{{\rm c}0}\!=\!0.0353$, there is a conspicuous peak in the background of $\Delta_{1}\!\sim\!0.17$ around the Fermi momentum $k\!=\!1$ at low Matsubara frequencies, which originates from the 1PR contribution $\Delta_{\rm c}^{\rm r}(\vec{k})\!\propto\!F(\vec{k})$.
Figure 5: The density of states around the excitation threshold calculated at five different temperatures.
The emergence of the 1PR structure in the self-energy necessarily accompanies a change in the single-particle density of states per volume
$D(\varepsilon)\!\equiv\!-\frac{2}{\pi V}\sum_{{\bf k}}{\rm Im}\,G(k,i\varepsilon_{n}\rightarrow\varepsilon\!+\!i0_{+}).$ (20)
Figure 5 plots the density of states over $-0.5\!\leq\!\varepsilon\!\leq 0.5$ at five different temperatures obtained by a numerical analytic continuation of the self-energy based on the 6-point Padé approximant VS77 . As seen there, the reduction in the density of states around the excitation threshold, called the pseudogap behavior, starts even from above $T_{1{\rm PR}}\!=\!0.100$ and become more and more conspicuous down through $T_{1{\rm PR}}$. The singular 1PR structure makes a reliable analytic continuation based on the 6-point Padé approximant impossible below $T\!=\!0.098$, which is an issue to be clarified in the future. The present mechanism for the pseudogap behavior may also be relevant to those observed in high-$T_{\rm c}$ superconductors, which still remains a matter of controversy TS99 ; NPK05 ; Matsuda17 . One of the possible candidates relevant to it in the present context is the electron-hole bubble $\chi_{GG}({\bf K}/2)$ at half the reciprocal lattice vector ${\bf K}$ corresponding to the antiferromagnetism. The scenario that $U\chi_{GG}({\bf K}/2)=1$ with $U>0$ is reached upon the emergence of the pseudogap without any explicit broken symmetries may be well worthy of giving detailed theoretical investigations.
It should be noted finally that the concept of the 1PR structure in the self-energy has emerged naturally in our effort of describing strongly correlated systems microscopically and numerically; without introducing Eqs. (9) and (14), we could not have obtained any finite solution of Dyson’s and the Dyson-Gor’kov equations below $T_{1{\rm PR}}$. In this respect, our model with a definite cutoff $k_{\rm c}\!=\!10$ may have been advantageous over the continuous one with ultraviolet divergences HRCZ07 in that no further renormalization of $U$, which may mask the singularities, is required.
IV Concluding Remarks
We have clarified existence of hidden singularities in correlated electron systems and developed a definite procedure to handle them. It is thereby shown that reaching them is characterized by the emergence of a novel structure in the self-energy called one-particle reducible (1PR), which necessarily accompanies a change in the single-particle density of states around the excitation threshold. The temperature at which it starts to develop can be regarded as a transition with no apparent broken symmetries.
The hidden singularities may have stood as a key untold difficulty in performing theoretical studies on correlated ordered phases based on the quantum field theory, as may be seen in the negligible number of theoretical studies on them compared with those on the normal state Yanase03 ; Kontani13 ; Scalapino12 , a resolution for which has been presented here. The 1PR structure has also been predicted to present in Bose-Einstein condensates based on a self-consistent $\Phi$ derivable scheme of satisfying Goldstone’s theorem and conservation laws Kita09 ; Kita14 . Thus, it may be a common feature over a wide range of correlated systems. Further detailed studies will be required to fully elucidate the novel phase with the 1PR structure in the self-energy.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP20K03848.
Appendix A Numerical Procedures
Numerical calculations of Eqs. (11) and (18) have been performed iteratively by choosing the noninteracting and mean-field Green’s functions as the starting points, respectively. Specifically for the normal state, we divide Green’s function in $\tau$ space with $\tau\in[0,\beta]$ as
$G(k,\tau)=G_{0}(k,\tau)+\delta G(k,\tau),$ (21)
where $G_{0}(k,\tau)$ is the noninteracting Green’s function
$G_{0}(k,\tau)=-\frac{e^{-\tau(k^{2}-\mu_{0})}}{e^{-\beta(k^{2}-\mu_{0})}+1},$ (22)
and $\delta G(k,\tau)$ denotes the correlation part updated iteratively starting from $\delta G(k,\tau)\!=\!0$. Variable $k$ is expressed as
$k=1+\sinh x,$ (23a)
and discretized with an equal interval in $x$ over $100\!\leq\!k\!\leq\!k_{\rm c}\!=\!$ 10; the number of points are $N_{k}\!=\!50\!\sim\!200$. Similarly, $\tau$ is expressed as
$\tau=\frac{\beta}{2}(1+\tanh x),$ (23b)
and discretized with an equal interval in $x$ over $\tau_{\rm c}\!\leq\!\tau\!\leq\!\beta\!-\!\tau_{\rm c}$ with $\tau_{\rm c}\!=\!10^{-8}$; the number of points are $N_{\tau}\!=\!50\!\sim\!400$.
The correlation functions of Eq. (4) are calculated in $\tau$ space based on the formula,
$\chi_{GG}(q,\tau)=\frac{1}{V}\sum_{\bf k}G(k,\tau)G(|{\bf k}-{\bf q}|,\beta-\tau),$ (24a)
$\chi_{G\bar{G}}(q,\tau)=-\frac{1}{V}\sum_{\bf k}G(k,\tau)G(|{\bf k}-{\bf q}|,\tau).$ (24b)
The angular integrals of $s\equiv{\bf k}\!\cdot\!{\bf q}/kq$ in Eq. (24) are performed over
${\max}\!\left(-1,\frac{k^{2}\!+\!q^{2}\!-\!k_{\rm c}^{2}}{2kq}\right)\!\leq s\leq 1,$ (25)
by dividing the interval at $(k^{2}\!+\!q^{2}\!-\!1)/2kq$ when it lies in it. The integration over each interval $s_{0}\!\leq\!s\!\leq s_{1}$ is performed by expressing
$s=\frac{s_{0}\!+\!s_{1}}{2}+\frac{s_{0}\!-\!s_{1}}{2}\tanh(\sinh x),$ (26)
and discretizing $x$ at an equal interval over $-2.5\!\leq\!x\!\leq\!2.5$; the number of integration points are $N_{s}\!=\!20\!\sim\!30$. The radial integrals of ${\bf k}$ in Eq. (24) are performed similarly over
${\rm max}(0,q-k_{\rm c})\leq k\leq k_{\rm c},$ (27)
by dividing the interval at the singular points $|1\!-\!q|$, $1$, $1\!+q$ when they lie in it, and making a change of variables similar to Eq. (26) for each interval.
Variable $q$ of Eq. (24) runs over $[0,2k_{\rm c}]$, which is divided at $q\!=\!2,k_{\rm c}$ into three regions. Each region is then discretized by the formula (26) with $s\rightarrow q$ with the relative number of points $4:3:3$; the total number of points is $N_{q}\!=\!50\!\sim\!200$.
The functions of Eq. (24) are then transformed into the $\omega_{\ell}$ space based on the method developed by Haussmann et al.HRCZ07 To be specific, we express any function $\chi(\tau)$ with $\tau\!\in\![\tau_{j},\tau_{j+1})$ by the cubic splineNRC
$\chi(\tau)\approx\chi_{j}(\tau)\equiv\sum_{n=0}^{3}a_{j,n}(\tau-\tau_{j})^{n},$ (28)
and carry out the Fourier transform by
[…]
$=$ $\,\sum_{n=0}^{3}a_{j,n}I_{j,n}(\omega_{\ell}),$ (29)
where $I_{j,n}(\omega_{\ell})$ is defined by
[…]
$=$ $\,e^{i\omega_{\ell}\tau_{j}}\sum_{\nu=0}^{n}\frac{n!}{\nu!}(-1)^{n-\nu}\frac{(\varDelta\tau_{j})^{\nu}\,e^{i\omega_{\ell}\varDelta\tau_{j}}-\delta_{\nu 0}}{(i\omega_{\ell})^{n-\nu+1}},$ (30)
with $\varDelta\tau_{j}\!\equiv\!\tau_{j+1}\!-\!\tau_{j}$. For $|\omega_{\ell}\varDelta\tau_{j}|\!<\!0.1$, this function is evaluated by expanding the formula in the Taylor series of $\omega_{\ell}\varDelta\tau_{j}$.
The boson Matsubara frequencies $\omega_{\ell}\!=\!2\pi\ell T$ ($\ell\!=\!0,1,2,\cdots$) for $\ell\!>\!\ell_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\omega_{\ell}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!\ell_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\omega_{\ell}\!\approx\!10^{7}$; the number of points are $N_{\omega}\!=\!50\sim 200$. Function $\bar{\chi}(\omega)$ with $\omega\!\in\![\omega_{\ell_{j}},\omega_{\ell_{j+1}})$ is then approximated by the cubic spline as
$\bar{\chi}(\omega)\approx$ $\,\bar{\chi}_{j}(\omega)\equiv\sum_{n=0}^{3}\bar{a}_{j,n}(\omega-\omega_{\ell_{j}}).$ (31)
The accuracy of Eq. (29) can be checked by the inverse transform
$\chi(\tau)=$ $\,\frac{1}{\beta}\sum_{\ell}\bar{\chi}(\omega_{\ell})\,e^{-i\omega_{\ell}\tau}$
$\approx$ $\,\frac{1}{\beta}\bar{\chi}(0)+2{\rm Re}\sum_{j=2}^{N_{\omega}-1}\bar{a}_{j,n}\bar{I}_{j,n}(\tau),$ (32)
where $\bar{I}_{j,n}(\tau)$ is defined by
$\bar{I}_{j,n}(\tau)\equiv$ $\,\frac{1}{\beta}\sum_{\ell=\ell_{j}}^{\ell_{j+1}-1}(\omega_{\ell}-\omega_{\ell_{j}})^{n}\,e^{-i\omega_{\ell}\tau}$
$=$ $\,\frac{e^{-i\omega_{\ell_{j}}\tau}}{\beta}\frac{d^{n}}{d(-i\tau)^{n}}\frac{e^{-i\varDelta\omega_{j}\tau}-1}{e^{-i\omega_{1}\tau}-1},$ (33a)
with $\varDelta\omega_{j}\!\equiv\!2\pi(\ell_{j+1}-\ell_{j})T$; the upper limit $\ell_{j+1}\!-\!1$ of Eq. (33a) for $j\!=\!N_{\omega}\!-\!1$ should be replaced by $\ell_{j+1}$. When $\varDelta\omega_{j}\tau<0.1$, we expand the last formula in $\varDelta\omega_{j}\tau$ as
$\bar{I}_{j,n}(\tau)=$ $\,\frac{e^{-i\omega_{\ell_{j}}\tau}}{\beta}\omega_{1}^{n}\sum_{\nu=0}^{\infty}(-1)^{\nu}\frac{(\tau\omega_{1})^{2\nu}}{(2\nu)!}\biggl{[}S_{n+2\nu}(L)$
$\,-\frac{i\tau\omega_{1}}{2\nu+1}S_{n+2\nu+1}(L)\biggr{]},$ (33b)
with $L\!\equiv\!\ell_{j+1}\!-\!\ell_{j}\!-\!1$ and
$S_{\nu}(L)\equiv\sum_{\ell=0}^{L}\ell^{\nu},$ (34)
and replace the upper limit $\infty$ in Eq. (33b) by $k_{\rm c}\!\sim\!10$ to obtain enough accuracy. The sum of Eq. (34) for $\nu\!\lesssim\!10$ can be evaluated by using Faulhaber’s formula. Equation (32) after the double Fourier transforms thereby obtained has been confirmed to reproduce Eq. (28) excellently.
Functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ calculated from Eq. (24) by Eq. (29) are then substituted in Eq. (5). The sum over $\ell$ is performed in $\tau$ space by using the Fourier transforms of
$U_{\rm ph}(\vec{q}\,)\equiv$ $\,U\Biggl{\{}U\chi_{GG}(\vec{q}\,)+\frac{3}{2}\,\frac{[U\chi_{GG}(\vec{q}\,)]^{2}}{1-U\chi_{GG}(\vec{q}\,)}$
$\,-\frac{1}{2}\,\frac{\bigl{[}U\chi_{GG}(\vec{q}\,)\bigr{]}^{2}}{1+U\chi_{GG}(\vec{q}\,)}\Biggr{\}},$ (35a)
$U_{\rm pp}(\vec{q}\,)\equiv$ $\,U\frac{[U\chi_{G\bar{G}}(\vec{q}\,)]^{2}}{1-U\chi_{G\bar{G}}(\vec{q}\,)},$ (35b)
as
$\Sigma_{\rm c}(k,\tau)=$ $\,\frac{1}{V}\sum_{{\bf q}}\bigl{[}U_{\rm ph}(q,\tau)G(|{\bf k}-{\bf q}|,\tau)$
$\,-U_{\rm pp}(q,\tau)G(|{\bf k}-{\bf q}|,\beta-\tau)\bigr{]},$ (36)
where the subscript of $\Sigma_{\rm c}$ denotes the correlation part. On the other hand, the angular and radial integrals of ${\bf q}$ in Eq. (36) are carried out in the same way as those over ${\bf k}$ in Eq. (24). Possible singular points in the radial integral over $q\!\in\![0,k\!+\!k_{\rm c}]$ are $|1\!-\!k|$, $1\!+\!k$, and $k_{\rm c}\!-\!k$.
Equation (36) thereby obtained is transformed into the $\varepsilon_{n}$ space in the same way as described from Eq. (32) through Eq. (34). The fermion Matsubara frequencies $\varepsilon_{n}\!=\!(2n\!+\!1)\pi T$ ($n\!=\!0,1,2,\cdots$) for $n\!>\!n_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\varepsilon_{n}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!n_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\varepsilon_{n}\!\approx\!10^{7}$; the number of points are $N_{\varepsilon}\!=\!50\sim 200$. Self-energy $\Sigma(\vec{k})$ thereby obtained is used to calculate $\delta G(\vec{k})\!\equiv\!G(\vec{k})\!-\!G_{0}(\vec{k})$, i.e., the correlation part of Green’s function. This $\delta G(\vec{k})$ is then transformed into the $\tau$ space by the procedure from Eq. (28) to Eq. (30) to obtain $\delta G(k,\tau)$ of the next iteration.
Below $T\!=\!T_{1{\rm PR}}$, we need to incorporate the second term of Eq. (10) to obtain any finite result. Calculations have been performed for several trial $\lambda_{\rm n}$’s to identify the value of $\lambda_{\rm n}$ at which $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ is met. It has turned out numerically that the value of $U\chi_{G\bar{G}}(\vec{0})$ changes almost linearly in terms of $\lambda_{\rm n}$ around $U\chi_{G\bar{G}}(\vec{0})\!=\!1$, so that the desired value of $\lambda_{\rm n}$ has been identified with no fundamental difficulties.
Similar calculations have been performed for the superconducting state iteratively starting from the mean-field solutions. The convergences are not bad at low temperatures of $T\!\lesssim\!0.2T_{{\rm c}0}$ even for the calculations of using Eq. (17). As we approach $T_{\rm c}\!\approx\!0.0624$, however, they become more and more unstable. Hence, we have set $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ from the beginning based on the results at low temperatures. This has enabled us to obtain the convergence near $T_{\rm c}$, which however has been reached only after many iterations.